Circular No. 7543

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2000ff, 2000fg, 2000fh, 2000fi, 2000fj, 2000fk, 2000fl Further to IAUC 7529, L.-G. Strolger reports the discovery of seven

additional apparent supernovae by the Nearby Galaxies Supernova Search team. Direct confirmations were made by Strolger and R. C. Smith (Kitt Peak 0.9-m telescope) on Nov. 11–19 and 22 and by P. Candia (Cerro Tololo 0.9-m telescope) on Nov. 18 and 19. Direct images obtained with various other telescopes indicate that SN 2000fg had brightened to R ∼ 18.2 on Nov. 28; that SN 2000fh appeared to be at or near maximum light (R ∼ 19.2) on Dec. 1; that SN 2000fk appeared to be just past maximum light at R ∼ 19.8 on Dec. 1; and that SN 2000fl appeared to be still rising at R ∼ 18.8 on Nov. 30.

SN 2000 UT α_{2000} δ_{2000} R Offset
2000ff Nov. 14 0^{h}45^{m}24.^{s}72 − 3^{o}16′47′′.2 19.5 3′′ E, 0′′.2 N
2000fg Nov. 16 1 24 31.36 −10 23 46.2 18.8 1′′.2 W, 0′′.8 S
2000fh Nov. 16 4 40 28.66 − 5 10 07.7 20.1 2′′.8 W, 0′′.3 N
2000fi Nov. 9 21 52 57.14 − 6 32 53.9 18.2 0′′.8 E, 2′′.1 N
2000fj Nov. 13 22 08 52.31 − 4 41 39.5 19.0 0′′.2 E, 2′′.0 S
2000fk Nov. 14 23 36 26.00 − 0 23 17.6 20.1 0′′.2 E, 1′′.6 N
2000fl Nov. 18 23 46 19.57 − 6 31 03.9 19.8 2′′.1 W, 1′′.4 S

R CORONAE BOREALIS J. A. Mattei, AAVSO, reports that this variable has begun to fade for

the first time since Aug. 1999–Feb. 2000, as indicated by the following visual magnitude estimates: Nov. 15.90 UT, 6.1 (G. Chaple, Jr., Townsend, MA); 16.73, 6.3 (G. Poyner, West Midlands, England); 22.73, 6.4 (R. J. Bouma, Groningen, The Netherlands); 25.70, 6.7 (S. Baroni, Milan, Italy); Dec. 1.20, 7.0 (T. Lange, Bovenden, Germany); 6.45, 7.0 (Chaple); 7.22, 7.6 (Bouma); 10.14, 8.1 (B. Andresen, Norway); 12.13, 9.0 (B. Granslo, Oslo, Norway); 13.45, 9.5 (Chaple); 14.70, 10.3 (Bouma).

COMET 41P/TUTTLE-GIACOBINI-KRESÁK This comet appears to be continuing its rapid brightening (cf. IAUC

7536), as indicated by the following m_{1} estimates: Dec. 5.82 UT, 11.1 (S. Yoshida, Ibaraki, Japan, 0.25-m reflector; visual); 6.52, 11.4 (A. Hale, Cloudcroft, New Mexico, 0.41-m reflector; visual); 15.83, 8.7 (K. Kadota, Ageo, Saitama, Japan, 0.18-m reflector + CCD; 3′ coma and 16′ tail).

2000 December 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green